a lab needs to make 100 gallons of a 18% acid solution by mixing a 12% acid solution with 20% solution. How many gallons of each solution is needed?
Answer by mananth(12269)

percent ---------------- quantity
12 ---------------- x
20 ---------------- 100-x
Total18.00%--------------100 ......
12x+20(100-x)=18*100
12x+2000 -20x=1800
12x-20x= 1800-2000
-8x=-200
/-8
x=25 liters of 12.00%
75 liters of 20.00%